Find the area of the portion of the plane 3x + 2y + z = 3 which lies above the annulus R given by 1 < x2 + y2 < 4.

Please show simple and clear steps for the problem above will leave a rating if so thank you
Answer #1

Area = \iint_{R} \sqrt{1+f_x ^2+f_y ^2} dA

where R is the region over which the function z=f(x,y) is required to find area.

Here, z = 3-3x-2y = f(x,y) , f_x = -3, f_y = -2

\sqrt{1+f_x^2+f_y^2} = \sqrt{14}

Area = \iint_{R}\sqrt{14}dA = \int_{0}^{2\pi}\int_{1}^{4} \sqrt{14} rdrd\theta = 2\pi\sqrt{14} \left [ \frac{r^2}{2} \right ]_1^4=15\sqrt{14} \pi